Previous

Calendar

October

2013

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

29

JD272

30

JD273

1

JD274

2

JD275

3

JD276
Site Visit (EG): Details , Photo (1)

open battery box?

4

JD277
at 2:56 MST, power outage at the site

at 3:34 MST, UPS battery exhausted, russter2 shutdown, data system off

at 4:15 MST, russter2 reboots

at around 4:19 MST, data system back up

First significant snowfall of the season in early morning (1-2"?).

5

JD278

6

JD279

7

JD280

8

JD281
Site Visit (SpB, ES, JT): Details , Photos (30)

Span cylinder at 1400 psi. N2 at 510 psi

from 9-10 MST moved EC-155 from 23m to 21.5m (located 40cm from LI-7200), moved LI-7500 (still at 21.5m)

at 12:55 MST, adjusted LI-6262 cal gas flow rate

at 13:00 MST, started ops66 (removed t.med, added co2.ec155), new FAST logger program (fast_18.csi).

turned off EC/Prof fans, powered up noah cr10x logger

moved three N2 and one span cylinder from MRS to C1

removed ati-k probe from the tower (in trailer)

CU Jeep

A few breezes in morning, nice in afternoon

9

JD282

10

JD283
Site Visit (EG, MR): Details , Photo (1)

EG checked battery box (it is dry)

check P.tot.18m (seems to have some dropouts)

11

JD284

12

JD285
Site Visit (MB): Details

teflon lines added to tower/picarro system

13

JD286

14

JD287
from: 10:05-10:32 MST, added the LI-7200 AIU-0366.metadata info

15

JD288
Snow

16

JD289

17

JD290

18

JD291

19

JD292

20

JD293

21

JD294

22

JD295
N2 is getting low!

23

JD296

24

JD297
Site Visit (SpB): Details , Photos (12+)

Span cylinder at 1300 psi. N2 at 0 psi --> 2150 psi

at 12:30 MST, swapped N2 cylinder; old=cc63143 new=cc329215

moved GPS power wire, uploaded new program to cr3000 (for cpec200)

started ops67 (made co2.ec155 hi-res, moved t.med to UC), new FAST logger program (fast_19.csi), new UC logger program (uc_04.csi).

at 15:00? MST, adjusted LI-6262 cal gas flow rate

at 13:20 MST, reduced LI-7200 inlet heater from 6W to 5W, uploaded new program to cr3000 (for cpec200)

powered up MRS cr10 logger

SpB Jeep

Sunny, calm. Some snow still on ground.

25

JD298
Site Visit (DS, EB): Details

work on molotch snow depth sensors

26

JD299

27

JD300

28

JD301
at around 10 MST, 2.5m LI-7500 went off-scale? (check this!)

29

JD302

30

JD303

31

JD304
Halloween
Windy (boo!)

1

JD305

2

JD306
Site Visit (SpB): Details , Photos (28)

Span cylinder at 1210 psi. N2 at 1950 psi

swapped LI-7200s (sn 72H-0479 replaced sn 72H-0192), LI-7550 box, and Fan Module (LI-7200-101)

reduced LI-7200 inlet heater from 5W to 4.6W,

couldn't connect to 2.5m LI-7500 (serial port problem?)

CU Jeep

Sunny and breezy. More snow on ground.